The Width of a rectangle is 5 meters less than its length. The area is 84 square meters. Find the length of a rectangle.
Answer by TimothyLamb(4379)

w = L - 5
Lw = 84
---
Lw = 84
L(L - 5) = 84
LL - 5L - 84 = 0
L^2 - 5L - 84 = 0
---
the above quadratic equation is in standard form, with a=1, b=-5, and c=-84
---
to solve the quadratic equation, by using the quadratic formula, plug this:
1 -5 -84
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
L = 12
L = -7
---
the negative root doesn't make sense for length, so use the positive root:
L = 12 meters
w = 7 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php